# Corrupted but Not Broken: Rethinking the Impact of Corrupted Data in Visual Instruction Tuning

### v1

@@ 7 Conclusions
In this paper, we show that while corrupted data hampers the performance of MLLMs, its impact is superficial. By disabling corruption-related parameters or post-training with clean data, the performance of MLLMs can be largely restored. Additionally, the corrupted MLLMs can effectively distinguish between clean and corrupted samples via self-validation, enabling self-cleaning of datasets. Building on this, the proposed corruption-robust training paradigm significantly outperforms existing strategies.
@@
Acknowledgment
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
@@ Appendix A Details in Figure 1
The “Simple FT” results in Figure 1 are aggregated from Figure 15 with LLaMA-3.1-8B (experiment details in Sec. 4.1). Results of “Ours” are aggregated from Figure 6 and 18 (experiment details in Sec. 6.1). The results of “Disable Params.” is taken from Table 1 with details in Sec. 4.2. The figure of precision under various corruption level is obtained from Figure 4 using LLaMA-3.1-8B at a recall of 0.5 (check Sec. 5.2 for more details).
@@ B.1 Corrupted Datasets
The prompt used by GPT (described in Sec 3) to generate corrupted samples is shown in Figure 7. We provide one example for each of the datasets used, an example of corrupted sample is shown in Figures 8-14. As can be seen, GPT produces corrupted samples in the following ways: changing the option letter in multiple choice VQA (Figure 9), replacing the correct answer in VQA with a plausible but incorrect one (Figures 10-13), inducing object hallucination in a conversation (Figure 8) and generating wrong captions (Figure 14).
@@ B.3 Evaluation Metrics
Appendix C More on Effect of Corruption
For experiments with uniform corruption, we vary the corruption ratio ($cr$) from 0% to 60% and construct a reference dataset with those corrupted samples removed to see whether they are contributing negatively. Figure 15 shows the model’s performance under uniform corruption across different ratios on various benchmarks.
As can be seen in Figure 15, for most tasks, corrupting the data results in worse performance than simply removing them. This degradation worsens with increasing $cr$, indicating the negative effect of corrupted data.
@@ D.2 Identifying Corruption-related Weights
Appendix E Details on Identifying Correctness using MLLM
E.1 Decision Rule and Evaluation
@@
E.2 Details on Scores
Perplexity
For a sample $x$, its perplexity is computed as :
The loss for this sample is $\log{\texttt{PPL}(x)}$.
E.3 Proof of Eq. (1)
By Bayes’ theorem,
$p(z=0|x_{c},\tilde{x}_{y})$ (3)
$=$ $\frac{p(\tilde{x}_{y}|z=0,x_{c})}{p(\tilde{x}_{y}|x_{c})}\cdot p(z=0|x_{c}).$
Since corruption is uniformly added on all samples, the correctness label $z$ is independent of $x_{c}$, and so $p(z=0|x_{c})=p(z=0)$. Substituting this into (3), we obtain
$p(z=1|x_{c},\tilde{x}_{y})$ $=1-c\cdot\frac{p(\tilde{x}_{y}|x_{c},z=0)}{p(\tilde{x}_{y}|x_{c})},$
where $c=p(z=0)$. Hence, we have
$p(z=1|x_{c},\tilde{x}_{y})-1\propto-\frac{p(\tilde{x}_{y}|x_{c},z=0)}{p(\tilde{x}_{y}|x_{c})}.$
E.4 Additional Results
Figure 16 shows the mean of $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ and $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ at varying corruption levels for clean samples. Compared to Figure 5 (Middle), a symmetrical but less pronounced trend is observed for clean samples.
@@ F.2 Sample Selection Methods
F.3 Scores Used for Sample Selection in Post-training
@@
When using these scores for sample selection in post-training, we experimented with both higher and lower values and found that selecting samples with lower values yielded better results (Table 9).
@@
Sample selection (post-training)
@@
Table 6: Comparisons of different corruption-robust strategies on Qwen-2.5-0.5B. Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
@@
Sample selection (post-training)
@@
Table 7: Comparisons of different corruption-robust strategies on Qwen-2.5-3B. Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
@@
Sample selection (post-training)
@@
Table 8: Comparisons of different corruption-robust strategies on Qwen-2.5-7B. Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
Figure 18: Average model performance of models post-trained (or fine-tuned from scratch) on data from data from different sources. Post-trained models are fine-tuned on dataset with various corruption levels (10%-40%) at first. Results with 50% corruption are in Figure 6.
@@
Table 9: Post-training on the bottom ($\downarrow$) and top ($\uparrow$) 30% subsets based on EL2N, GradNorm, and Entropy scores. The reported results represent the average performance. All models are initially fine-tuned on data with a corruption ratio of $cr=50\%$.

### v2

@@ 6 Conclusions
In summary, we show that the negative impact of corrupted data on MLLMs is largely reversible and that these models can inherently identify corrupted samples. Leveraging these insights, we introduce a corruption-robust training method that outperforms existing strategies, improving the robustness of Visual Instruction Tuning.
@@
Ethic Statement
There is no ethical problem in our study.
@@ Appendix A Details in Figure 2
The “Simple FT” results in Figure 2 are aggregated from Figure 16 with LLaMA-3.1-8B (experiment details in Sec. 4). Results of “Ours” are aggregated from Figure 7 and 19 (experiment details in Sec. 5.2). The results of “Disable Params.” is taken from Table 1 with details in Sec. 4.2. The figure of precision under various corruption level is obtained from Figure 5(b) using LLaMA-3.1-8B at a recall of 0.5 (check Sec. 5.1 for more details).
@@ B.1 Corrupted Datasets
The prompt used by GPT-4o (described in Sec 3) to generate corrupted samples is shown in Figure 8. We provide examples of corrupted samples for each of the datasets used in Figures 9-15. As can be seen, GPT produces corrupted samples in the following ways: changing the option letter in multiple choice VQA (Figure 10), replacing the correct answer in VQA with a plausible but incorrect one (Figures 11-14), inducing object hallucination in a conversation (Figure 9) and generating wrong captions (Figure 15).
@@ B.3 Evaluation Metrics
Appendix C More on Effects of Corrupted Data
Due to limited space, Figure 3 only presents a subset of evaluation tasks used in this paper. Here we provide the results for all tasks in Figure 16.
@@ D.2 Identifying Corruption-related Weights
Appendix E Details on Identifying Correct Samples using MLLM
E.1 Details on Scores
Perplexity
For a sample $x$, its perplexity is computed as :
The loss for this sample is $\log{\texttt{PPL}(x)}$.
E.2 Decision Rule and Evaluation
@@
E.3 Analysis on Improved Understanding of Corrupted Samples
To understand why improved understanding of corrupted samples lead to the distribution shift of $\hat{p}(z=1|x_{c},\tilde{x}_{y};\bm{\theta})$ on corrupted samples, we start by rewriting $p(z=1\mid x_{c},\tilde{x}_{y})$ using Bayes’ rule as
$p(z=1\mid x_{c},\tilde{x}_{y})=1-p(z=0\mid x_{c},\tilde{x}_{y}).$ (2)
By Bayes’ theorem, we have
$p(z=0\mid x_{c},\tilde{x}_{y})$
$=$ $\frac{p(\tilde{x}_{y}\mid z=0,x_{c})}{p(\tilde{x}_{y}\mid x_{c})}\cdot p(z=0\mid x_{c}).$ (3)
Since corruption is uniformly applied to all samples, the correctness label $z$ is independent of $x_{c}$, and thus
$p(z=0\mid x_{c})=p(z=0).$
Substituting this into (2) and (3), we get
$p(z=1\mid x_{c},\tilde{x}_{y})=1-c\cdot\frac{p(\tilde{x}_{y}\mid x_{c},z=0)}{p(\tilde{x}_{y}\mid x_{c})},$
where $c=p(z=0)$. Therefore,
$p(z=1\mid x_{c},\tilde{x}_{y})-1\propto-\frac{p(\tilde{x}_{y}\mid x_{c},z=0)}{p(\tilde{x}_{y}\mid x_{c})}.$
This shows that $p(z=1|x_{c},\tilde{x}_{y})-1$ is proportional to the negative of the ratio between $p(\tilde{x}_{y}|x_{c},z=0)$ and $p(\tilde{x}_{y}|x_{c})$. Note $p(\tilde{x}_{y}|x_{c},z=0)$ can be regarded as the MLLM’s understanding of corrupted samples, which we estimate through explicitly prompting the MLLM similar to Val_PPL as follows:
[…]
Figure 17 shows the means of the estimated probabilities $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ and $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ at varying corruption levels over the corrupted datasets for fine-tuning. We observe a sharp rise in $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ (blue) when transitioning from a clean model to one trained with slight corruption ($cr=10\%$), followed by a slower growth as the corruption increases further. In contrast, $\hat{p}(\tilde{x}_{y}|x_{c};\bm{\theta})$ (orange) exhibits a modest increase and remains consistently lower than the other probability.
This shows the MLLM’s understanding of corrupted $\hat{p}(\tilde{x}_{y}|x_{c},z=0;\bm{\theta})$ gets significantly improved when only including 10% of corruptions.
@@ F.2 Sample Selection Methods
F.3 Scores Used for Sample Selection in Further Fine-tuning
@@
When using these scores for sample selection in further fine-tuning, we experimented with both higher and lower values and found that selecting samples with lower values yielded better results (Table 9).
@@
Further fine-tuning
@@
Table 6: Comparisons of different corruption-robust strategies at a corruption ratio of 50% on LLaVA-1.5 (Qwen-2.5-0.5B). Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
@@
Further fine-tuning
@@
Table 7: Comparisons of different corruption-robust strategies at a corruption ratio of 50% on LLaVA-1.5 (Qwen-2.5-3B). Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
@@
Further fine-tuning
@@
Table 8: Comparisons of different corruption-robust strategies at a corruption ratio of 50% on LLaVA-1.5 (Qwen-2.5-7B). Here, Avg. refers to the average performance on 11 benchmarks (normalized to 0-100). Best results are Bold, second best are underlined.
Figure 19: Average model performance of models further fine-tuned (or fine-tuned from scratch) on data from data from different sources. Further fine-tuned models are trained on dataset with various corruption levels (10%-40%) at first. Results with 20% and 50% corruption are in Figure 7.
@@
Table 9: Further fine-tuning on the bottom ($\downarrow$) and top ($\uparrow$) 30% subsets based on EL2N, GradNorm, and Entropy scores. The reported results represent the average performance. All models are initially fine-tuned on data with a corruption ratio of $cr=50\%$.